OpenStudy (anonymous):

Find the LCD of the set of fractions: 7/x^2+4x+4 and 5/4-x^2

OpenStudy (loser66):

where are you stuck?

OpenStudy (danjs):

\[\frac{ 7 }{ x^2 } + \frac{ 4x }{ 1 } + \frac{ 4 }{ 1 }\] and \[\frac{ 5 }{ 4 } - \frac{ x^2 }{ 1 }\]

OpenStudy (danjs):

actually i am going to assume you have parenthesis in there

OpenStudy (danjs):

\[\frac{ 7 }{ x^2+4x+4 } ~~~and~~~ \frac{ 5 }{ 4-x^2 }\]

OpenStudy (danjs):

Factoring each results in- \[\frac{ 7 }{ (x+2)(x+2) }~~~and~~~\frac{ 5 }{ -(x-2)(x+2) }\]

OpenStudy (danjs):

The LCD is : (x - 2)( x + 2)(x + 2) for \[\frac{ 7(x-2) }{ (x-2)(x+2)(x+2) }~~and~~\frac{ -5(x+2) }{ (x-2)(x+2)(x+2) }\]